OpenStudy (anonymous):

integrate (cos 2x- cos 2 b) / (cos x - cos b)

OpenStudy (noelgreco):

With respect to what?

OpenStudy (anonymous):

x

OpenStudy (turingtest):

that is\[\int{\cos(2x)-\cos(2b)\over\cos x-\cos b}dx\]? because if those 2's were exponents this would be a lot easier

OpenStudy (turingtest):

@Aparna2 I can't help you if you don't answer me

hartnn (hartnn):

write cos2x as \(2cos^2 x -1\) cos 2b as \(2cos^2 b-1\) then u convert \(cos 2x-cos 2b \) as \(2(cos^2 x-cos^2 b)\) now u can write that as (cos x +cos b)(cos x -cos b) and cancel out cosx -cos b

OpenStudy (turingtest):

holy heck, so it turns out the same as if the 2's were exponents! mind blown...

OpenStudy (turingtest):

...except there's a 2 in front of the integral now

hartnn (hartnn):

yup. to check: the final answer would be 2(x cos b + sin x ) +c